Eco Plastics Company Since its inception, Eco Plastics Company has been revolutionizing plastic and trying to do its part to save the environment. Eco biodegradable plastic that her company is marketing to manufacturing companies throughout the southeastern United States. After operating as a private company for years, Eco went public in 2015 and it is listed on the Nasdaq stock exchange. s founder, Marion Cosby, developed a 6 As the chief financial officer of a young company with lots of investment opportunities, Ecos CFO closely monitors the firms cost of capital. The CFO keeps tabs on each of the individual costs of Ecos three main financing sources: long-term debt, preferred and common stock. The target capital structure for Eco is given by the weights in the following table Sources of capital Weight Long-term debt Preferred stock Common stock equity | 50% Total 30% 20% 100% At the present time, Eco can raise debt by selling 20-year bonds with a $1,000 par value and a 11% annual coupon interest rate. Ecos corporate tax rate is 40%, and its bonds generally require an average discount of $40 per bond and flotation costs of $35 per bond when being sold. Ecos outstanding preferred stock pays a 9% dividend and has a $100 per share par value. The cost of issuing and selling additional preferred stock is expected to be $10 per share. Because Eco is a young firm that requires lots of cash to grow it does not currently pay a dividend to common stockholders. To track the cost of common stock the CFO uses the capital asset pricing model (CAPM). The CFO and the firms investment advisors believe that the appropriate risk-free rate is 4% and that the markets expected return equals 12%. Using data from 2015 through 2018, Ecos CFO estimates the firms beta to be 1.5. Although Ecos current target capital structure includes 20% preferred stock, the company is considering using debt financing to retire the outstanding preferred stock, thus shifting their target capital structure to 50% long-term debt and 50% common stock. If Eco shifts its capital mix from preferred stock to debt, its financial advisors expect its beta to increase to 1.7.
A. Calculate Ecos current after-tax cost of long-term debt. B. Calculate Ecos current cost of preferred stock.
C. Calculate Ecos current cost of common stock. D. Calculate Ecos current weighted average cost of capital.
Answer #1
A) The issue price per bond = 1000-40-35 = $           925
YTM (using an online calculator), with price
of $925, n = 20, C = 11% (annual coupon) = 12.00%
After tax cost of debt = 12%*(1-40%) = 7.20%
B) Current cost of preferred stock = 9/(100-10) = 10.00%
C) Current cost of common stock (CAPM) = risk free rate+beta*(expected market retutn-risk free rate) = 4%+1.5*(12%-4%) = 16.00%
D) Current WACC = 7.20%*30%+10%*20%+16%*50% = 12.16%